OpenStudy (anonymous):

@math&ing001 help in trigoo plzzzz

OpenStudy (anonymous):

dude...can u help me in trignometry..plzz teach me how to do it

OpenStudy (math&ing001):

What do you need help with ?

OpenStudy (anonymous):

all of trignometry..i s am weak in trigg

OpenStudy (anonymous):

plzzz

OpenStudy (math&ing001):

I need to know where you're at in trig so I can help. Did you see sin, cos and tan, the unit circle ? Maybe just post an exercise you're having difficulty with and we can go from there.

OpenStudy (anonymous):

i dont know how to prove l.h.s = r.h.s i tried a lot...i dont know what to do..i am even ready to give my both kidneys away if thats what is required to know how to prove as those type of quetions are asked in threefourth of the paper

OpenStudy (math&ing001):

lhs = rhs is the form of any equation, that doesn't need proof.

OpenStudy (anonymous):

no in trignometry how do we solve themm

OpenStudy (anonymous):

plzz teach meee

OpenStudy (anonymous):

@ganeshie8 plz helpp

ganeshie8 (ganeshie8):

try khan https://www.khanacademy.org/math/trigonometry

ganeshie8 (ganeshie8):

understanding SOH-CAH-TOA and unit circle are the most import things when you review trig

OpenStudy (anonymous):

okieeee

OpenStudy (anonymous):

no like iwwnat to prove l.h.s =r.h.s using identtiess

ganeshie8 (ganeshie8):

for example ?

OpenStudy (math&ing001):

Here's a list of identities you might need. http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

okieee

OpenStudy (anonymous):

lemme chekkk

ganeshie8 (ganeshie8):

good luck!

OpenStudy (anonymous):

i will give u an example...

OpenStudy (anonymous):

prove that tan thitha + cot thitha =2/x^2-1

OpenStudy (math&ing001):

I don't know what you're trying to prove here since you have two diffrent variables, but if you're looking to find x as function of theta : \[\tan \theta + \cot \theta = \frac{ \sin^2 \theta + \cos^2 \theta }{ \sin \theta \cos \theta }=\frac{ 2 }{ \sin 2 \theta }=\frac{ 2 }{ \cos(2 \theta - \pi / 2)}=\frac{ 2 }{ [\sqrt{2}\cos( \theta - \pi / 4)]^{2}-1 }\]